# FleSpeech: Flexibly Controllable Speech Generation with Various Prompts

### v1

@@ 3.4 Training Strategy
In the first stage, the model is trained on a combination of two datasets to achieve basic speech synthesis capabilities with the large-scale corpus ensuring stability. In the second stage, the model is fine-tuned on high-expressive data to achieve domain alignment. In the third stage, we freeze the generation model backbone and start training the multimodal encoder to enable the model to support modal inputs other than speech prompts. Notably, during this stage, the multimodal prompt encoder is updated with the generation loss in addition to the diffusion loss.
@@ 4 Multimodal Dataset
Due to the scarcity of multimodal controllable speech synthesis data, we propose a method for constructing such a database. Compared to existing data, the collected data is not only larger in scale but also includes facial modality with richer facial annotation information. Details about the collected data and comparisons with other multimodal speech synthesis datasets can be found in Appendix B.
@@ 5.2 Evaluation Metrics
Objective metrics includes Word Error Rate (WER), Speaker Similarity (SPK-Sim), UTMOS Saeki et al. (2022)  [^7] https://github.com/tarepan/SpeechMOS, Emotion Accuracy, Gender Accuracy, and other speech attribute accuracy. Details about these objective metrics can be found in Appendix C.1.
Subjective metrics include the Mean Opinion Score (MOS) to evaluate speech naturalness (N-MOS) and similarity (Sim-MOS). Higher N-MOS means better naturalness while higher Sim-MOS indicates better similarity with the specific target. Details about the subjective metrics can be found in Appendix C.2
@@ 6 Experimental Results
We evaluated FleSpeech using both single-type prompts and various combinations of prompt types. Additionally, the extended capabilities of FleSpeech, including speech editing and voice conversion, were also assessed. The introduction to the various comparison methods, including MM-TTS Guan et al. (2024), Salle Ji et al. (2024a), NaturalSpeech2 Shen et al. (2024), and PromptTTS2 Leng et al. (2024) can be found in the Appendix D.
@@ Appendix A Model Configurations
The language model for semantic prediction adopts the LLaMA architecture with 16 layers and 16 attention heads. The hidden size and intermediate size are 1024 and 4096, respectively. The flow matching for the acoustic feature prediction is based on the DiT architecture with 12 layers, 128 attention heads, and a hidden dimension of 768. For MPE, the number of queries in QueryMLP is set to 16, with 6 layers, 6 attention heads, and an intermediate size of 256. The reference audio encoder consists of 6 attention blocks with a hidden size of 512. During both the training and inference stages, the length of the audio prompt is fixed at 6 seconds.
Appendix B Details of Collected Data
@@
Appendix C Evalution Metrics
C.1 Objective Metrics
@@
For the accuracy of other speech attributes, we utilise the previously mentioned pipeline for style label annotation to extract attribute values and compare their relative magnitudes across different labels. For example, the speech rate associated with the "fast speaking rate" label exceeds that of the "slow speaking rate" label. For face attribute evaluation, we extract speaker embeddings from MPE and use a face classifier  [^10] https://github.com/lsimmons2/bmi-project to predict Body Mass Index (BMI). Additionally, we apply the DeepFace Serengil and Ozpinar (2021) model to determine gender, race, and age. We then train an MLP-based predictor to infer facial attributes from the speaker embeddings, comparing the predicted attributes against the provided facial descriptions to compute the accuracy.
C.2 Subjective Metrics
In the subjective evaluation, each sample was rated on a scale from 1 to 5, with increments of 0.5, based on its similarity to the reference utterance. The score of MOS test ranges from 1 to 5 with an interval of 0.5, in which 1 means very bad and 5 means excellent. Both N-MOS and Sim-MOS are reported with a 95% confidence interval. We select 50 speech samples for each test, which are listened by 20 listeners for subjective evaluations.
@@ C.2 Subjective Metrics
Appendix D Comparison models

### v2

@@ 3.4 Training Strategy
In the first stage, the model is trained on a combination of two datasets to achieve basic speech synthesis capabilities with the large-scale corpus ensuring stability. In the second stage, the model is fine-tuned on high-expressive data to achieve domain alignment. In the third stage, we freeze the generation model backbone and start training the multimodal encoder to enable the model to support modal inputs other than speech prompts. Notably, during this stage, the multimodal prompt encoder is updated with the generation loss in addition to the diffusion loss. The details of the training objective can be found in Appendix B.
@@ 4 Multimodal Dataset
Due to the scarcity of multimodal controllable speech synthesis data, we propose a method for constructing such a database. Compared to existing data, the collected data is not only larger in scale but also includes facial modality with richer facial annotation information. Details about the collected data and comparisons with other multimodal speech synthesis datasets can be found in Appendix C.
@@ 5.2 Evaluation Metrics
Objective metrics includes Word Error Rate (WER), Speaker Similarity (SPK-Sim), UTMOS Saeki et al. (2022)  [^7] https://github.com/tarepan/SpeechMOS, Emotion Accuracy, Gender Accuracy, and other speech attribute accuracy. Details about these objective metrics can be found in Appendix D.1.
Subjective metrics include the Mean Opinion Score (MOS) to evaluate speech naturalness (N-MOS) and similarity (Sim-MOS). Higher N-MOS means better naturalness while higher Sim-MOS indicates better similarity with the specific target. Details about the subjective metrics can be found in Appendix D.2
@@ 6 Experimental Results
We evaluated FleSpeech using both single-type prompts and various combinations of prompt types. Additionally, the extended capabilities of FleSpeech, including speech editing and voice conversion, were also assessed. The introduction to the various comparison methods, including MM-TTS Guan et al. (2024), Salle Ji et al. (2024a), NaturalSpeech2 Shen et al. (2024), and PromptTTS2 Leng et al. (2024) can be found in the Appendix E.
@@ Appendix A Model Configurations
The language model for semantic prediction adopts the LLaMA architecture with 16 layers and 16 attention heads. The hidden size and intermediate size are 1024 and 4096, respectively. The flow matching for the acoustic feature prediction is based on the DiT architecture with 12 layers, 12 attention heads, and a hidden dimension of 768. For MPE, the number of queries in QueryMLP is set to 16, with 6 layers, 6 attention heads, and an intermediate size of 256. The reference audio encoder consists of 6 attention blocks with a hidden size of 512. During both the training and inference stages, the length of the audio prompt is fixed at 6 seconds.
The diffusion model consists of a diffusion process and a denoising process. For the diffusion process, given the audio embedding $x_{0}$, the forward diffusion process transforms it into Gaussian noise under the noise schedule $\beta$ as follows:
$\mathrm{d}x_{t}=-\frac{1}{2}\beta_{t}x_{t}\mathrm{t}+\sqrt{\beta_{t}}\mathrm{d}\omega_{t},t\in[0,1]$ (1)
For the denoising process, the denoising process aims to transform the noisy representation $x_{t}$ to the audio embedding $x_{0}$ by the following formulation Song et al. (2021).
$\mathrm{d}x_{t}=-\frac{1}{2}\left(x_{t}+\nabla\log p_{t}\left(x_{t}\right)\right)\beta_{t}\mathrm{~{}d}t,\quad t\in[0,1]$ (2)
The diffusion module is trained to estimate the gradients of log-density of noisy data ($\bigtriangledown logp_{t}(z_{t})$) by predicting the origin audio embedding $x_{0}$, conditioned on the embeddings from different prompt modalities, noised audio embedding, and diffusion step $t$ that indicates the degree of noise in the diffusion model.
@@
Appendix B Training Objective
In the first stage, the language model performs the next token prediction task and is optimized using the cross-entropy loss. Meanwhile, flow matching reconstructs the hidden layer features and is optimized with $L_{2}$ loss. In the second stage, the MPE is optimized using the $L_{1}$ loss calculated between the output embedding $Pred\ x_{0}$ derived from different prompt modalities and ground-truth embedding $x_{0}$ obtained from the audio modality, in addition to the loss function of the first stage.
To achieve flexible control, the MPE applies masking to the received prompts of different modalities. Specifically, the audio modality prompt, serving as the target for the MPE, remains consistently present. For data containing both text and audio modality prompts, the MPE maps the text prompt embeddings to the audio embeddings without any masking. In the case of data that includes prompts from all three modalities, there is a one-third probability of masking the text style description, a one-third probability of masking the facial image and facial description, and a one-third probability of not applying any masking. This strategy enables the model to accept various combinations of prompts as input during the inference stage.
Appendix C Details of Collected Data
Appendix D Evalution Metrics
D.1 Objective Metrics
@@
For the accuracy of other speech attributes, we utilize the previously mentioned pipeline for style label annotation to extract attribute values and compare their relative magnitudes across different labels. For example, the speech rate associated with the “fast speaking rate” label exceeds that of the “slow speaking rate” label. For face attribute evaluation, we extract speaker embeddings from MPE and use a face classifier  [^10] https://github.com/lsimmons2/bmi-project to predict Body Mass Index (BMI). Additionally, we apply the DeepFace Serengil and Ozpinar (2021) model to determine gender, race, and age. We then train an MLP-based predictor to infer facial attributes from the speaker embeddings, comparing the predicted attributes against the provided facial descriptions to compute the accuracy.
D.2 Subjective Metrics
In the subjective evaluation, each sample was rated on a scale from 1 to 5, with increments of 0.5 based on its similarity to the reference utterance, where a score of 1 indicates “very bad” and a score of 5 signifies “excellent.” Both Normalized Mean Opinion Score (N-MOS) and Similarity Mean Opinion Score (Sim-MOS) are reported with a 95% confidence interval. We selected 50 speech samples for each test, which were listened to by 20 listeners for subjective evaluations.
@@ D.2 Subjective Metrics
Appendix E Comparison models
@@
Appendix F Visualizing the Relationship between Facial Attributes and Voice
Figure 4: Fundamental Frequency (F0) curve of the speech at different ages and BMI levels groups by gender
To further validate that FleSpeech can establish associations between facial attributes and voice characteristics, we extracted the fundamental frequency (F0) from synthesized speech prompted by different BMI and age groups. As shown in the upper two panels of Fig. 4, the F0 of older women decreases as age increases. In contrast, for elderly males, despite vocal cord atrophy, the F0 tends to increase. The lower two panels of Fig. 4 reveal that being overweight tends to cause articulation difficulties, leading to a decrease in F0 for females, whereas males experience an increase in F0. These findings align with conclusions from prior research, indicating that the proposed FleSpeech can capture variations in speech characteristics across different ages and BMI levels.
Appendix G Visualization of MPE Embedding Space
We design MPE to encode prompts from different modalities into a unified space. To validate this, we utilized the MPE to extract embeddings corresponding to each single-modality prompt. Considering that the MPEs in the language model and flow matching do not share parameters, we conducted analyses on both. The test set comprised 2000 randomly selected sentences containing prompts from all three modalities, including 20 speakers with 200 sentences each. The MPE outputs are projected to 2D by t-SNE Van der Maaten and Hinton (2008). Each color represents a modality.
As illustrated in Fig. 5, both MPEs exhibited similar trends: embeddings mapped by the MPE from different modalities reside within the same embedding space and are not partitioned into multiple subspaces where partitioning into subspaces would imply that each modality is encoded separately, failing to capture the intermodal relationships. Furthermore, the embeddings from audio prompts demonstrated stronger clustering, indicating that audio prompts are more directional than text and facial prompts. In contrast, text and facial prompts exhibit a one-to-many relationship with voice attributes, showing more significant variability.
(a) MPE in language model
(b) MPE in flow matching
Figure 5: TSNE visualization of MPE output embedding clustering.